# Dynamics of the sine-Gordon equation on tadpole graphs

### v1

@@ A.2 Morse index estimate
Let $A$ be a densely defined lower semi-bounded symmetric operator (that is, $A\geq mI$) with finite deficiency indices, $n_{\pm}(A)=k<\infty$, in the Hilbert space ${H}$, and let $\widehat{A}$ be a self-adjoint extension of $A$. Then the spectrum of $\widehat{A}$ in $(-\infty,m)$ is discrete and consists of, at most, $k$ eigenvalues counting multiplicities.
Next, let $\mathcal{G}_{0}$ be the graph determined by the set $\mathbb{E}=(-L,0)\cup(0,L)\cup(L,+\infty)$ and $H^{2}(\mathcal{G}_{0})=H^{2}(-L,0)\oplus H^{2}(0,L)\oplus H^{2}(L,+\infty)$. For $\mathbb{U}=(u,w,v)\in H^{2}(\mathcal{G}_{0})$ we define the Laplacian operator $-\Delta_{0}$ on $\mathcal{G}_{0}$ in the natural form
$-\Delta_{0}\mathbb{U}=(-u^{\prime\prime},-w^{\prime\prime},-v^{\prime\prime}).$ (A.9)
With this notations we have the following results (via a translation, all the statements below can be considered for $\mathbb{E}_{1}=(-L,0)\cup(0,L)\cup(0,+\infty)$ with vertex of the graph at $\nu=0$).
Theorem A.6.
The Schrödinger type operator $-\Delta_{0}$ in (A.9) on $L^{2}(\mathcal{G}_{0})$, with domain
$D(-\Delta_{0})=\{(u,w,v)\in H^{2}(\mathcal{G}_{0}):$ $u(-L)=w(L)=v(L)=0,\;u(0-)=w(0+)=0,$ (A.10)
$u^{\prime}(0-)=w^{\prime}(0+),\;w^{\prime}(L-)-u^{\prime}(-L+)=v^{\prime}(L+)\},$
is a densely defined symmetric operator with deficiency indices $n_{\pm}(-\Delta_{0})=2$. Therefore, we obtain that $(-\Delta_{0},D(-\Delta_{0}))$ has a four-parameter family of self-adjoint extensions. In particular, we have the family a two-parameter $(-\Delta_{0},D_{\lambda,\gamma})_{(\lambda,\gamma)\in\mathbb{R}^{2}}$ with
$D_{\lambda,\gamma}=\{(u,w,v)\in H^{2}(\mathcal{G}_{0}):$ $\;u(0-)=w(0+),\;u(-L)=w(L)=v(L),\;w^{\prime}(0+)-u^{\prime}(0-)=\lambda w(0+),$ (A.11)
$w^{\prime}(L-)-u^{\prime}(-L+)=v^{\prime}(L+)+\gamma v(L)\}.$
Thus, for $\lambda=0$, the family $(-\Delta_{0},D_{0,\gamma})_{(0,\gamma)\in\mathbb{R}^{2}}$ can be identified with the family $(-\Delta,D_{\gamma})_{\gamma\in\mathbb{R}}$ defined in Theorem A.4. In other words, $(-\Delta,D_{\gamma})_{\gamma\in\mathbb{R}}$ belongs to the self-adjoint extensions of the symmetric operator $(-\Delta_{0},D(-\Delta_{0}))$.
Proof.
The symmetry property of $(-\Delta_{0},D(-\Delta_{0}))$ is immediate. The density property of $D(-\Delta_{0})$ in $L^{2}(\mathcal{G}_{0})$ follows from $C_{c}^{\infty}(-L,0)\bigoplus C_{c}^{\infty}(0,L)\bigoplus C_{c}^{\infty}(L,+\infty)\subset D(-\Delta_{0})$. Now, by following a similar analysis as in Proposition A.5 in [14], the adjoint operator $(-\Delta_{0}^{*},D(-\Delta_{0}^{*}))$ of $(-\Delta_{0},D(-\Delta_{0}))$ is given by
$-\Delta_{0}^{*}=-\Delta_{0},\quad D(-\Delta_{0}^{*})=\{(u,w,v)\in H^{2}(\mathcal{G}_{0}):u(0-)=w(0+),\;u(-L)=w(L)=v(L)\}.$ (A.12)
Next, the deficiency subspaces $\mathcal{D}_{\pm}=\ker(-\Delta_{0}^{*}\mp i)$ have dimension 2. Indeed, by using the definition of $D(-\Delta_{0}^{*})$ in (A.12) we can see that $\mathcal{D}_{\pm}=\mathrm{span}\{\Pi_{\pm i},\Theta_{\pm i}\}$ with
$\Pi_{-i}=\Big{(}\underset{\bf-L<x<0}{e^{\sqrt{i}(x+L)}},\underset{\bf 0<x<L}{e^{-\sqrt{i}(x-L)}},\underset{\bf L<x}{e^{\sqrt{i}(x-L)}}\Big{)},\;\;\Theta_{-i}=\Big{(}\underset{\bf-L<x<0}{e^{-\sqrt{i}(x+L)}},\underset{\bf 0<x<L}{e^{\sqrt{i}(x-L)}},\underset{\bf L<x}{e^{\sqrt{i}(x-L)}}\Big{)},\;\;\mathrm{Im}(\sqrt{i})<0,$
and
$\Pi_{+i}=\Big{(}\underset{\bf-L<x<0}{e^{\sqrt{-i}(x+L)}},\underset{\bf 0<x<L}{e^{-\sqrt{-i}(x-L)}},\underset{\bf L<x}{e^{\sqrt{-i}(x-L)}}\Big{)},\;\;\Theta_{+i}=\Big{(}\underset{\bf-L<x<0}{e^{-\sqrt{-i}(x+L)}},\underset{\bf 0<x<L}{e^{\sqrt{-i}(x-L)}},\underset{\bf L<x}{e^{\sqrt{-i}(x-L)}}\Big{)},\;\;\mathrm{Im}(\sqrt{-i})>0.$
Thus, from the Krein and von Neumann theory, we deduce that $(-\Delta_{0},D(-\Delta_{0}))$ has a four-parameter family of self-adjoint extensions, in particular, we obtain one family a two-parameter of self-adjoint extensions $(-\Delta_{0},D_{\lambda,\gamma})_{(\lambda,\gamma)\in\mathbb{R}^{2}}$ determined by $D_{\lambda,\gamma}$ in (A.11). The statement about $\lambda=0$ is immediate. This finishes the proof. ∎
The domains $D_{\lambda,\gamma}$ can be seen as $\delta$-interaction type domains with “deltas” located at the points $x=0,L$. The introduction of the graph $\mathcal{G}_{0}$ is a main piece for obtaining a Morse index estimate of $(\mathcal{L}_{Z},D_{Z})$ in (5.1) (see Theorem 5.6 and Theorem A.7 below). We note that this strategy of splitting the graph $\mathcal{G}$ at $x=0$ is due to the fact that the positive single-lobe profile $\phi_{1}$ on $[-L,L]$ satisfies $\phi_{1}^{\prime}(0)=0$ and so Proposition A.5 can be used.
Theorem A.7.
Let $(\phi_{1},\phi_{2})\in D_{Z}$ be a positive single-lobe kink state for system (1.6). Then, for $(\mathcal{L}_{Z},D_{Z})$ defined in (5.1) we have the Morse index estimate, $n(\mathcal{L}_{Z})\leqq 2$.
Proof.
The idea of the proof is to split the operator $\mathcal{L}_{Z}$ on a convenient graph and to apply the extension theory of symmetric operators. Without loss of generality, we consider $c_{1}=c_{2}=1$. Then, for the metric graph $\mathcal{G}_{0}$ defined above, we consider the diagonal-matrix $\mathcal{M}=\mathrm{diag}(M_{i})_{1\leqq i\leqq 3}$ with
$M_{1}u(x)$ $=(-\partial_{x}^{2}+\cos(\phi_{1}))\,u(x),$ $x\in(-L,0)$ (A.13)
$M_{2}w(x)$ $=(-\partial_{x}^{2}+\cos(\phi_{1}))\,w(x),$ $x\in(0,L)$
$M_{3}v(x)$ $=(-\partial_{x}^{2}+\cos(\phi_{2}))\,v(x),$ $x>L.$
Then, $(\mathcal{M},D(\mathcal{M}))$, with $D(\mathcal{M})=D(-\Delta_{0})$ in (A.10) is a symmetric operator and the deficiency indices satisfy $n_{\pm}(\mathcal{M})=n_{\pm}(-\Delta_{0})=2$ (because $\mathcal{M}-\Delta_{0}$ is a bounded operator on $L^{2}(\mathcal{G}_{0})$). Moreover, there is a two-parameter family of self-adjoint extensions $(\mathcal{M}_{\lambda,\gamma},D(\mathcal{M}_{\lambda,\gamma}))_{(\lambda,\gamma)\in\mathbb{R}^{2}}$ associated to $(\mathcal{M},D(\mathcal{M}))$ with $\mathcal{M}_{\lambda,\gamma}=\mathcal{M}$ and $D(\mathcal{M}_{\lambda,\gamma})=D_{\lambda,\gamma}$ defined in (A.11). Next, we see $\mathcal{M}\geqq 0$. Indeed, for ${\bf{U}}=(u,w,v)\in D(-\Delta_{0})$ and from the (well-defined) formulae
$M_{1}u$ $=-\frac{1}{\phi_{1}^{\prime}}\frac{d}{dx}\Big{[}(\phi_{1}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{u}{\phi_{1}^{\prime}}\Big{)}\Big{]},\quad$ $x\in(-L,0),$ (A.14)
$M_{2}w$ $=-\frac{1}{\phi_{1}^{\prime}}\frac{d}{dx}\Big{[}(\phi_{1}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{w}{\phi_{1}^{\prime}}\Big{)}\Big{]},\quad$ $x\in(0,L),$
$M_{3}v$ $=-\frac{1}{\phi_{2}^{\prime}}\frac{d}{dx}\Big{[}(\phi_{2}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{v}{\phi_{2}^{\prime}}\Big{)}\Big{]},\quad$ $x>L,$
we obtain via integration by parts the relation
[…]
where the $I_{i}$’s are the remainders defined by
$I_{1}$ $=-\frac{u}{\phi_{1}^{\prime}}(\phi_{1}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{u}{\phi_{1}^{\prime}}\Big{)}\Big{|}_{-L}^{0},$ (A.16)
$I_{2}$ $=-\frac{w}{\phi_{1}^{\prime}}(\phi_{1}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{w}{\phi_{1}^{\prime}}\Big{)}\Big{|}_{0}^{L},$
$I_{3}$ $=-\frac{v}{\phi_{2}^{\prime}}(\phi_{2}^{\prime})^{2}\frac{d}{dx}\Big{(}\frac{v}{\phi_{2}^{\prime}}\Big{)}\Big{|}_{L}^{\infty}.$
Next, we see $I_{i}=0$ for $i=1,2,3$. Indeed, since $u(0-)=u(-L)=0$, $\phi_{1}^{\prime}(-L)\neq 0$, $\phi_{1}^{\prime}(0)=0$ and $\phi_{1}^{\prime\prime}(0)\neq 0$, it follows from L’Hopital’s rule that
$I_{1}=\phi_{1}^{\prime\prime}(0)\lim_{\epsilon\to 0^{-}}\frac{u^{2}(\epsilon)}{\phi_{1}^{\prime}(\epsilon)}=0.$
Similarly, we get $I_{2}=0$. For $I_{3}$ we have from $v(L)=0$ and $\phi_{2}^{\prime}(L)\neq 0$,
$I_{3}=v(L)v^{\prime}(L)-v^{2}(L)\frac{\phi_{2}^{\prime\prime}(L)}{\phi_{2}^{\prime}(L)}=0.$
Due to Proposition A.5 and Theorem A.6, the family of self-adjoint extensions $(\mathcal{M}_{\lambda,\gamma},D(\mathcal{M}_{\lambda,\gamma}))_{(\lambda,\gamma)\in\mathbb{R}^{2}}$ satisfies $n(\mathcal{M}_{\lambda,\gamma})\leqq 2$ for every $(\lambda,\gamma)\in\mathbb{R}^{2}$, and for $\lambda=0$ we get $n(\mathcal{L}_{Z})\leqq 2$. This finishes the proof. ∎
A.3 Perron-Frobenius property
@@
In what follows, we establish the Perron-Frobenius property for the family of self-adjoint operators (~$(\widetilde{\mathcal{L}},D_{\delta,\gamma})$, where
~$\widetilde{\mathcal{L}}=-c_{2}^{2}\partial_{x}^{2}+\cos(\psi_{\mathrm{even}}),$ (A.17)
@@
for any $\gamma\in\mathbb{R}$. So, we start with the following remark: by Weyl’s essential spectrum theorem (cf. Reed and Simon [45]), we have that the essential spectrum, $\sigma_{\mathrm{ess}}(\widetilde{\mathcal{L}})$, of $\widetilde{\mathcal{L}}$ satisfies $\sigma_{\mathrm{ess}}(\widetilde{\mathcal{L}})=[1,+\infty)$ for any $\gamma$.
@@
Consider the family of self-adjoint operators defined in (A.17)-(A.18), (~$(\widetilde{\mathcal{L}},D_{\delta,\gamma})_{\gamma\in\mathbb{R}}$. For $\gamma$ fixed, assume that $\beta=\inf\sigma(\widetilde{\mathcal{L}})<1$ is the smallest eigenvalue. Then, $\beta$ is simple, and its corresponding eigenfunction $\zeta_{\beta}$ is positive (after replacing $\zeta_{\beta}$ by $-\zeta_{\beta}$ if necessary) and even.
@@
This result follows by a slight twist of standard abstract Perron-Frobenius arguments (see Proposition 2 in Albert et al. [3]). Without loss of generality we consider $c_{2}=1$. The basic point in the analysis is to show that the Laplacian operator $-\Delta_{\gamma}\equiv-\frac{d^{2}}{dx^{2}}$ on the domain $D_{\delta,\gamma}$ has its resolvent $R_{\mu}=(-\Delta_{\gamma}+\mu)^{-1}$ represented by a positive kernel for some $\mu>0$ sufficiently large. Namely, for $f\in L^{2}(\mathbb{R})$
[…]
with $K(x,y)>0$ for all $x,y\in\mathbb{R}$. By convenience of the reader we show this main point, the remainder of the proof follows the same strategy as in [3]. Thus, for $\gamma$ fixed, let $\mu>0$ be sufficiently large (with $-2\sqrt{\mu}<\gamma$ in the case $\gamma<0$), then from the Krein formula (see Theorem 3.1.2 in [4]) we obtain
$K(x,y)=\frac{1}{2\sqrt{\mu}}\Big{[}e^{-\sqrt{\mu}|x-y|}-\frac{\gamma}{\gamma+2\sqrt{\mu}}e^{-\sqrt{\mu}(|x|+|y|)}\Big{]}.$
Moreover, for every $x$ fixed, $K(x,\cdot)\in L^{2}(\mathbb{R})$. Thus, the existence of the integral above is guaranteed by Holder’s inequality. Now, since $K(x,y)=K(y,x)$, it is sufficient to show that $K(x,y)>0$ in the following cases.
  1. (1)
    Let $x>0$ and $y>0$ or $x<0$ and $y<0$: for $\gamma\geqq 0$, we obtain from $\frac{\gamma}{\gamma+2\sqrt{\mu}}<1$ and $|x-y|\leqq|x|+|y|$, that $K(x,y)>0$. For $\gamma<0$ and $-2\sqrt{\mu}<\gamma$, follows immediately $K(x,y)>0$.
  2. (2)
    Let $x>0$ and $y<0$: in this case,
    $K(x,y)=\frac{1}{\gamma+2\sqrt{\mu}}e^{-\sqrt{\mu}(x-y)}>0$
    for any value of $\gamma$ (where again $-2\sqrt{\mu}<\gamma$ in the case $\gamma<0$) .

### v2

@@ A.1 Extension theory for the Laplacian operator on a tadpole graph
Let $A$ be a densely defined lower semi-bounded symmetric operator (that is, $A\geq mI$) with finite deficiency indices, $n_{\pm}(A)=k<\infty$, in the Hilbert space H{H}, and let $\widehat{A}$ be a self-adjoint extension of $A$. Then the spectrum of $\widehat{A}$ in $(-\infty,m)$ is discrete and consists of, at most, $k$ eigenvalues counting multiplicities.
A.2 Perron-Frobenius property
@@
In what follows, we establish the Perron-Frobenius property for the family of self-adjoint operators $(\widetilde{\mathcal{L}},D_{\delta,\gamma})$, where
$\widetilde{\mathcal{L}}=-c_{2}^{2}\partial_{x}^{2}+\cos(\psi_{\mathrm{even}}),$ (A.3)
@@
[…]
@@
Consider the family of self-adjoint operators defined in (A.3)-(A.4), $(\widetilde{\mathcal{L}},D_{\delta,\gamma})_{\gamma\in\mathbb{R}}$. For $\gamma$ fixed, assume that $\beta=\inf\sigma(\widetilde{\mathcal{L}})<1$ is the smallest eigenvalue. Then, $\beta$ is simple, and its corresponding eigenfunction $\zeta_{\beta}$ is positive (after replacing $\zeta_{\beta}$ by $\zeta_{\beta}$ if necessary) and even.
@@
This result follows by a slight twist of standard abstract Perron-Frobenius arguments (see Proposition 2 in Albert et al. [3]) applied to Schrödinger operators with point interactions. We recommend the reader to see the proof of Theorem 6.7 in Angulo [5]. ∎
A.3 Proof of Lemma 4.10
Proof.
In the sequel we determine the correct values of $k$ and $Z$ such that $(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$, namely, such that we have the relation
$2\phi^{\prime}_{1}(L)=\phi^{\prime}_{2}(L)+Z\phi_{2}(L),$
with $\phi_{1}=\phi_{1,k}$, $\phi_{2}=\phi_{2,a(k)}$, and $a(k)$ satisfying relation in (4.13). We divide this study into several cases.
$\bullet$ Case (1). Let $\frac{L}{c_{1}}>\frac{\pi}{2}$. Consider the unique $k_{0}\in(0,1)$ such that $K(k_{0})=\frac{L}{c_{1}}$. Then for $k>k_{0}$ it follows that $K(k)>\frac{L}{c_{1}}$. Now, from properties of the Jacobi elliptic functions (see [19]), we obtain from (4.13) (the continuity condition at the vertex $\nu=L$) the following limits:
$\lim_{k\to 1}\operatorname{sech}\Big(\frac{a(k)}{c_{2}}\Big)$ $=\lim_{k\to 1}\frac{k^{\prime}}{\mathrm{dn}(L/c_{1};k)}=\frac{0}{\operatorname{sech}(L/c_{1})}=0,$ (A.5)
$\frac{a(k)}{c_{2}}=-\ln\Big(\frac{\mathrm{dn}(L/{c_{1}};k)+k\mathrm{cn}(L/{c_{1}};k)}{k^{\prime}}\Big)\equiv j(k).$ (A.6)
Then, numerical simulations always show that $j(k)$ in (A.6) is strictly decreasing for $k\in(k_{0},1)$ and so $a^{\prime}(k)<0$, for $k\in(k_{0},1)$ (see Figure 16 for the graph of $j=j(k)$, $k\in(k_{0},1)$ with $c_{1}=1$, $L=\pi$ and $k_{0}\approx 0.984432$). We note that by using formula (710.53) in [19] and deriving with regard to $k$ both sides of equality in (4.13) we can also see that $a(k)$ is strictly decreasing after long calculations.
Figure 16: Graph of the function $j=j(k)$ in (A.6) for $k\in(k_{0},1)$, with $c_{1}=1$, $L=\pi$ and where $k_{0}\approx 0.984432$.
$\phi^{\prime}_{1}(L)=\frac{2k}{c_{1}}\mathrm{cn}(L/c_{1}+K;k)=-\frac{2kk^{\prime}\operatorname{sn}(L/c_{1};k)}{c_{1}\mathrm{dn}(L/c_{1};k)}.$ (A.7)
From (1.8) and (4.13) we get for $\phi_{2}=\phi_{2,a}$,
$\phi^{\prime}_{2}(L)=-\frac{2}{c_{2}}\operatorname{sech}\Big(\frac{a}{c_{2}}\Big)=-\frac{2}{c_{2}}\frac{k^{\prime}}{\mathrm{dn}(L/c_{1};k)}.$ (A.8)
$\frac{k^{\prime}}{c_{1}\mathrm{dn}(L/c_{1};k)}\Big[\frac{c_{1}}{2c_{2}}-k\operatorname{sn}(L/c_{1};k)\Big]=Z\beta,$ (A.9)
$Z=\frac{\operatorname{sech}(a/c_{2})}{c_{1}\arctan(e^{-a/c_{2}})}\Big[\frac{c_{1}}{2c_{2}}-k\operatorname{sn}(L/c_{1};k)\Big]\equiv H(k).$ (A.10)
$H(k)=Z,$ (A.11)
for $Z\in(-\infty,\frac{2}{\pi c_{2}})$ fixed, and some $k\in(k_{0},1)$ with $a=a(k)$ solving (4.13). In other words, we will apply the Intermediate Value Theorem. Next, we consider the following cases:
  1. (I)
    Let $Z>0$. Then, a priori, $k$ needs to satisfy $k\operatorname{sn}(L/c_{1};k)<\frac{c_{1}}{2c_{2}}$ for solving (A.11). Thus, from basic properties of the Jacobi elliptic functions we get the following limits
    $\lim_{k\to 1}k\operatorname{sn}(L/c_{1};k)$ $=\tanh(L/c_{1}),$ (A.12)
    Moreover, the mapping
    $k\in(k_{0},1)\to g(k)\equiv k\operatorname{sn}(L/c_{1};k),$ (A.13)
    is strictly increasing (we note that for $k<k_{0}$ it is possible to have $g(k)\leqq 0$ and with $g$ having some oscillations; see Figures 17 and 18), and so we get
    Figure 17: Graph of the function $g(k)$ defined on (A.13) for $k\in(0,1)$, with $c_{1}=1$, $L=2.5\pi$.
    Let us consider the following subcases:
    1. (i)
      Suppose $\frac{c_{1}}{2c_{2}}\geqq\tanh(L/c_{1})$. Then $H(k)>0$ for $k\in(k_{0},1)$. Now, from (A.5) and properties of the Jacobi elliptic functions (cf. [19]), we obtain
      $\lim_{k\to 1}H(k)$ $=\frac{0}{\arctan(\infty)}\Big[\frac{c_{1}}{2c_{2}}-\tanh(L/c_{1})\Big]=0,$ (A.14)
    2. (ii)
      Suppose $k_{0}<\frac{c_{1}}{2c_{2}}<\tanh(L/c_{1})$. Then, there is a unique $\gamma\in(k_{0},1)$ with $g(\gamma)=\frac{c_{1}}{2c_{2}}$ and such that for $k\in(k_{0},\gamma)$ follows $H(k)>0$. From $\lim_{k\to\beta}H(k)=0$ and (A.14), we have for $Z\in\big(0,\frac{4}{\pi c_{1}}\big[\frac{c_{1}}{2c_{2}}-k_{0}\big]\big)$ that there is a unique $k_{Z}\in(k_{0},\beta)$ with $H(k_{Z})=Z$. This shows $(1)(ii)$.
  2. (II)
    Now let $Z<0$. From (A.10) follows that $k$ a priori needs to satisfy $\frac{c_{1}}{2c_{2}}<k\operatorname{sn}(L/c_{1};k)<\tanh(L/c_{1})$. Let us examine the following subcases:
    1. (iii)
    2. (iv)
    3. (v)
  3. (III)
    Now suppose that $Z=0$. From (A.10) we need to have $H(k)=0$ and so $\frac{c_{1}}{2c_{2}}=k\operatorname{sn}(L/c_{1};k)$ is true if and only if $\frac{c_{1}}{2c_{2}}\in(k_{0},\tanh(L/c_{1}))$ and in this case there is a unique $k\in(k_{0},1)$ such that $H(k)=0$. This shows $(1)(vi)$.
$\bullet$ Case (2). Now consider the case $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$. Then $K(k)>\frac{L}{c_{1}}$ for all $k\in(0,1)$. Thus, since $\mathrm{dn}(\cdot;0)=1$ we get from (4.13) and (A.5) again $\lim_{k\to 1}a(k)=$-$\infty$ and $\lim_{k\to 0}a(k)=0$. Moreover, the mapping $k\in(0,1)\to g(k)=k\operatorname{sn}(L/c_{1};k)$ is strictly increasing (see Figure 18) with $k\operatorname{sn}(L/c_{1};k)>0$, $\lim_{k\to 1}k\operatorname{sn}(L/c_{1};k)=\tanh(\frac{L}{c_{1}})$ and $\lim_{k\to 0}k\operatorname{sn}(L/c_{1};k)=0$.
Once again, we split the analysis into subcases:
  1. (I)
    Let $Z>0$. Then we have:
    1. (i)
    2. (ii)
    This proves $(2)(i)$.
  2. (II)
  3. (III)
  1. (i)
  2. (ii)
  1. (i)
  2. ii)
The Lemma is now proved. ∎